Use the formula A=P(l=r)t to find the rater r at which $3000 grows to $5247 in 4 years
Answer by stanbon(75887)

Use the formula A=P(l+r)t to find the rate r at which $3000
grows to $5247 in 4 years
------
5247 = 3000*(1+r)^4
-----
1.749 = (1+r)^4
------
(1+r) = 1.749^(1/4)
----
r = 1.1500 - 1
----
r = 0.15 = 15%
=================
Cheers,
Stan H.
-------------------